OpenStudy (anonymous):

The area of a rectangle is 21ft^2 , and the length of the rectangle is 1ft less than twice the width. Find the dimensions of the rectangle.

OpenStudy (abmon98):

A=L*W L=2W-1 W=W A=(2W-1)W 21=2W^2-W 2W^2-W-21=0 Solve Quadratic equation.